The average fuel consumption of vehicles driving the roads can be modeled by
F(t) = 0.425t2 − 16.442t + 840.325
where F(t) represents the gallons per vehicle per year t years since 1970.
(a) What was the average fuel consumption in 1975? (Round your answer to two decimal places. Include units with your numerical answers.)
According to the model, in 1975 the average fuel consumption per vehicle per year was______.
(b) What was the average fuel consumption in 1985? (Round your answer to two decimal places. Include units with your numerical answers.)
According to the model, in 1985 the average fuel consumption per vehicle per year was_______.
(c) Find the vertex for this model and explain its meaning. (Round your answer to two decimal places. Include units with your numerical answers.)
The vertex is (t, F) (______), which means that in______the average fuel consumption per vehicle per year was_______.

Answer by stanbon(75887)

The average fuel consumption of vehicles driving the roads can be modeled by
F(t) = 0.425t2 − 16.442t + 840.325
where F(t) represents the gallons per vehicle per year t years since 1970.
(a) What was the average fuel consumption in 1975? (Round your answer to two decimal places. Include units with your numerical answers.)
According to the model, in 1975 the average fuel consumption per vehicle per year was F(5) = 0.425*5^2 - 16.442^5 + 840.325
------------------
(b) What was the average fuel consumption in 1985? (Round your answer to two decimal places. Include units with your numerical answers.)
According to the model, in 1985 the average fuel consumption per vehicle per year was F(10) = 0.425*10^2 - 16.442*10 + 840.325
------------------------
(c) Find the vertex for this model and explain its meaning. (Round your answer to two decimal places. Include units with your numerical answers.)
Vertex occurs when t = -b/(2a) = 16.442/(2*0.425) = 19.34
The corresponding F(t) value is F(19.34)
---------------------------
Cheers,
Stan H.
--------------